If f(x) = -2x+5 and g(x)=x-4X=1, find each value.

Answer by tutorcecilia(2152)

f(x) = -2x+5
f(1) = -2(1)+5
f(1) = -2+5
f(1) = 3
.
g(x)= x-4
g(1) = (1)-4
g(1) = -3